OpenStudy (anonymous):

A golf ball is dropped from a height of 81 inches. It rebounds to 2/3 of its original height and continues rebounding in this manner. How far does it travel before coming to rest?

OpenStudy (mathmale):

Sounds like a problem involving geometric series. I'd suggest you draw a diagram showing each drop and each rebound. If this were just a simple geometric series, you'd write \[\sum_{n=1}^{infinity}a*r ^{n-1}, n: {1,2,3,...}\] but there's more to the present problem than that. What;'s extra?

OpenStudy (anonymous):

when the golf ball is first dropped it will cover 81 m then it rebounds to cover 2/3 of 81 =54m then it falls back 54m then it rebounds 2/3 of 54 =36m falls back 36m so total distance = 81 +(54+54+36+36+.....+0 ) 81 +2(an G.P with first term = 54 comm diff = 2/3) 81 +2(162) (USING INFINITE G.P FORMULA ) =405m i found this online but what do they mean by ( an G.P with first term)

OpenStudy (mathmale):

What do y ou mean by G. P.? I think the G stands for Geometric; the P possibly for Progression.

OpenStudy (anonymous):

im not sure what It means but I think G is for geometric and p for series

OpenStudy (rock_mit182):

GEOmetric progression

OpenStudy (rock_mit182):

also known as geometric sequence

OpenStudy (rock_mit182):

\[a _{n}=a _{n-1}*r\] for every integer n >= 1.

OpenStudy (rock_mit182):

or also \[a _{n+1} = a _{n} *r\]

OpenStudy (rock_mit182):

where r ≠ 0 is the common ratio and a is a scale factor, equal to the sequence's start value